# VideoLifter: Lifting Videos to 3D with Fast Hierarchical Stereo Alignment

### v2

@@ 3.3 Hierarchical Gaussian Alignment
In this stage, we perform dense 3D scene reconstruction using Gaussian splitting. First, we construct local 3D Gaussians within each fragment, and then merge these local models via hierarchical Gaussian alignment. The key design question is how to construct a globally coherent 3D scene while preserving local scene details without incurring significant alignment errors.
@@
Guassian initialization: In the local fragment registration step, we obtain the key frame’s dense point cloud, along with the relative poses and scale factors for the other frames, which can be used to obtain entire point map within fragment. To initialize $G^{f}_{i}$, we then assign a Gaussian to each point in the pixel-wise point cloud, setting its attributes as: color based on the corresponding pixel, center at the 3D point location, opacity adhering to the 3D-GS protocol [21], and scaling such that it projects as a one-pixel radius in the 2D image (by dividing the depth by the focal length). We set Gaussians as isotropic to reduce the degrees of freedom in Gaussian training.
Further refinement: The initial camera poses and point cloud positions may contain minor inaccuracies, we further refine them through joint optimization of camera poses and Gaussian parameters. Specifically, for each local Gaussian in $G^{f}_{i}$, we randomly sample frames within the fragment, render the current Gaussians into sampled frame, and backpropagate gradient updates to the Gaussian positions, colors, scales, opacities, and camera poses.
@@
Hierarchical Gaussian Alignment:
Next, we merge the local fragment-level 3D Gaussian sets $\mathcal{G}^{f}=\{G^{f}_{i}\}_{i=1}^{m}$ to the final consistent 3D scene. Naive pairwise progressive merging poses Challenge ➀: an excessive number of Gaussians for optimization, and Challenge ➁: inconsistencies among the various local Gaussian sets. To avoid these issues, we propose a tree-based hierarchical pipeline (see Fig. 4) that iteratively performs three key processes.
2) Pair-wise merging with visibility-mask-driven Gaussian pruning (Fig. 4 yellow): To avoid duplicating Gaussians in regions where $G^{f}_{1}$ already provides adequate scene reconstruction, we use a visibility mask to determine areas that $G^{f}_{1}$ could faithfully reconstruct:
$\small M(\mathbf{p})$ $=\text{Conf}(\mathbf{p})>\beta+D(\mathbf{p})>0,$ (2)
where $\mathbf{p}$ represents pixel position on image plane, $D(\mathbf{p})$ is rendered depth. $\text{Conf}(\mathbf{p})=\sum_{i}\alpha_{i}(\mathbf{p})\prod_{j=1}^{i-1}\left(1-\alpha_{j}(\mathbf{p})\right)$ is the rendered confidence, e.g. if a pixel contains information from the current gaussian, and threshold $\beta$ determines the masking criteria.
As in Fig. 5, the visibility mask effectively identifies regions where Gaussians from $G^{f}_{1}$ could provide sufficient depth and confidence. Importantly, the mask remains robust to occlusions, even when covered by large Gaussians. As our initialization is pixel-wise point clouds, the inverse of the visibility mask can be directly applied to $G^{f}_{2}$ to select Gaussians that complement the missing regions of $G^{f}_{1}$. These selected Gaussians inherit previously optimized parameters, ensuring seamless integration into the unified representation.
@@ 4.1 Experimental Setup
Datasets: We carry out comprehensive experiments on various real-world datasets, including Tanks and Temples [22], CO3D-V2 [35]. For Tanks and Temples, following NoPe-NeRF [6] and CF-3DGS [14], we assess both novel view synthesis quality and pose estimation accuracy across 8 scenes, spanning both indoor and outdoor environments. In each case, we use 7 out of every 8 frames from the video clips as training data and evaluate the novel view synthesis on the remaining frame except Family. For CO3D-V2, containing thousands of object-centric videos where the camera orbits the object, recovering camera poses is significantly more challenging due to the complex and large camera motions. We follow the protocol of CF-3DGS [14] to select the same 10 scenes from different object categories and apply the same procedure to divide training and testing sets.
@@
NeRF-mm [46] - $\sim$13h33min 0.5313 20.02 0.5450 0.035
BARF [24] GT K $\sim$20h 0.6075 23.42 0.5362 0.078
NoPe-NeRF [6] GT K $\sim$30h 0.7125 25.49 0.4113 0.013
CF-3DGS [14] GT K $\sim$2h20min 0.9213 31.14 0.0859 0.004
@@
Metrics: We assess our approach on two key tasks: generating novel views and estimating camera poses. For the task of novel view synthesis, we evaluate performance using common metrics such as PSNR, SSIM [45], and LPIPS [49]. In terms of camera pose estimation, we evaluate Absolute Trajectory Error (ATE) [42] and utilize COLMAP-generated poses from all dense views as our ground-truth for evaluation. While Relative Pose Error (RPE) [42] evaluates the local consistency of relative transformations between consecutive frames, it can be sensitive to discrepancies in intrinsic parameters such as focal length. However, ATE provides a more comprehensive measure of global trajectory accuracy and is better aligned with the goals of our method, which emphasizes globally consistent 3D reconstruction [42]. As such, we prioritize ATE as the primary metric for evaluating the poses of VideoLifter.
@@ 4.2 Quantitative Evaluations
To quantitatively evaluate the quality of synthesized novel views, we present the results in Tab. 3 for the Tanks and Temples dataset and Tab. 4 for the CO3D-V2 dataset. Baseline models were re-trained using their officially released code to ensure a fair comparison of training time. Compared to other self-calibrating radiance field methods, our approach achieves superior performance in terms of efficiency and rendering quality, which is largely thanks to our decoupled fragment registration and hierarchical alignment process. Compared to the most relevant baseline CF-3DGS [14], we reduce $>$80% training time yet get $>$0.012 LPIPS improvement on Tanks and Temples, and reduce $>$85% training time yet get $>$0.12 LPIPS improvement on CO3D-V2 dataset. Note that our VideoLifter does not require any ground-truth camera parameters, making it more adaptable to scenes that do not have or fail to get precomputed intrinsics from COLMAP. Compared to NeRFmm [46], which also does not need ground-truth camera parameters, our VideoLifter delivers much better quality and much less training time. Detailed per-scene breakdown results could be found in the Supplementary.
@@
NeRF-mm [46] - $\sim$17h22min 0.4380 13.43 0.7058 0.061
NoPe-NeRF [6] GT K $\sim$35h 0.7030 25.54 0.5190 0.055
CF-3DGS [14] GT K $\sim$2h55min 0.6821 22.98 0.3515 0.014
@@ 4.3 Qualitative Evaluations
As shown in Fig. 6, for large-scale scenes from the Tanks and Temples dataset, thanks to the hierarchical design in VideoLifter, our method consistently produces sharper details among all test views, and preserves fine details that are well-optimized within each fragment. For the CO3D-V2 dataset, which includes 360-degree scenes with complex trajectories, achieving a globally consistent 3D reconstruction without any COLMAP initialization is even more challenging. Baselines that rely on monocular depth prediction to unproject images into point clouds often suffer from depth scale inconsistencies, making them fragile and prone to failure. Even CF-3DGS, which uses the more robust ZoeDepth monocular depth estimator [5], encounters severe failures on CO3D-V2. In contrast, VideoLifter leverages 3D geometry priors to achieve robust registration, making it highly adaptable and resilient in challenging settings. More results could be found in the Supplementary.
@@
Local: Use MASt3R MVS Init. [12] 35min 0.8582 27.91 0.1768
@@
4.4 Ablation Study
Local Fragment Registration. Replacing it with direct MASt3R multi-view stereo initialization increases training time and lowers reconstruction quality, suggesting that outputs directly from MASt3R is less accurate and introduces errors to gaussian optimization, especially in the challenging long video sequences.
Hierarchical Gaussian Alignment. Removing our hierarchical alignment and instead adding local Gaussians sequentially (as in CF-3DGS [14]) prolongs training and degrades performance, underscoring the efficiency and accuracy gains from our hierarchical design.
@@ 4.4 Ablation Study
Confidence Threshold $\beta$. Setting $\beta$ too low allows fewer Gaussians to merge—leading to under-reconstructed areas—while a high $\beta$ merges too many Gaussians, complicating and slowing down training.
5 Conclusion
We introduced VideoLifter, a framework that efficiently reconstructs 3D scenes from monocular video without requiring pre-computed camera poses or predefined intrinsics. VideoLifter leverages learning-based stereo priors for initial sparse scene reconstruction and incorporates a hierarchical alignment approach with 3D Gaussian splatting to produce dense, globally consistent models. Compared to previous state-of-the-art methods [14, 6], VideoLifter achieves high-quality reconstructions from casual video, reducing computational demand.

### v3

@@ 3.3 Hierarchical Gaussian Alignment
In this stage, we perform dense 3D scene reconstruction using Gaussian Splatting. First, we construct local 3D Gaussians within each fragment, and then merge these local models via hierarchical Gaussian alignment. The key design question is how to construct a globally coherent 3D scene while preserving local scene details without incurring significant alignment errors.
@@
Guassian initialization: In the local fragment registration step, we obtain the key frame’s dense point cloud, along with the relative poses and scale factors for the other frames, which can be used to obtain entire point map within fragment. To initialize $G^{f}_{i}$, we then assign a Gaussian to each point in the pixel-wise point cloud, setting its attributes as: color based on the corresponding pixel, center at the 3D point location, opacity adhering to the 3D-GS protocol kerbl20233d , and scaling such that it projects as a one-pixel radius in the 2D image (by dividing the depth by the focal length). We set Gaussians as isotropic to reduce the degrees of freedom in Gaussian training.
Further refinement: The initial camera poses and point cloud positions may contain minor
@@
inaccuracies, we further refine them through joint optimization of camera poses and Gaussian parameters. Specifically, for each local Gaussian in $G^{f}_{i}$, we randomly sample frames within the fragment, render the current Gaussians into sampled frame, and backpropagate gradient updates to the Gaussian positions, colors, scales, opacities, and camera poses.
Hierarchical Gaussian Alignment:
Next, we merge the local fragment-level 3D Gaussian sets $\mathcal{G}^{f}=\{G^{f}_{i}\}_{i=1}^{m}$ to the final consistent 3D scene. Naive pairwise progressive merging poses Challenge ➀: an excessive number of Gaussians for optimization and ➁: inconsistencies among various local Gaussian sets. To avoid these issues, we propose a tree-based hierarchical pipeline ( Fig. 4) that iteratively performs three key processes.
2) Pair-wise merging with visibility-mask-driven Gaussian pruning (Fig. 4 yellow): To avoid duplicating Gaussians in regions where $G^{f}_{1}$ already provides adequate scene reconstruction, with $\mathbf{p}$ (the pixel position on image plane), we use a visibility mask to determine areas that $G^{f}_{1}$ could faithfully reconstruct: $M(\mathbf{p})=\text{Conf}(\mathbf{p})>\beta+D(\mathbf{p})>0.$ where $D(\mathbf{p})$ is rendered depth. $\text{Conf}(\mathbf{p})=\sum_{i}\alpha_{i}(\mathbf{p})\prod_{j=1}^{i-1}\left(1-\alpha_{j}(\mathbf{p})\right)$ is the rendered confidence, e.g. if a pixel contains information from the current gaussian, and threshold $\beta$ determines the masking criteria.
@@
Fig. 5 shows that the visibility mask effectively identifies regions where Gaussians from $G^{f}_{1}$ could provide sufficient depth and confidence. Importantly, the mask remains robust to occlusions, even when covered by large Gaussians. As our initialization is pixel-wise point clouds, the inverse of the visibility mask can be directly applied to $G^{f}_{2}$ to select Gaussians that complement the missing regions of $G^{f}_{1}$. These selected Gaussians inherit previously optimized parameters, ensuring seamless integration into a unified representation.
@@ 4.1 Experimental Setup
Datasets: We carry out comprehensive experiments on various real-world datasets, including Tanks and Temples knapitsch2017tanks , CO3D-V2 reizenstein2021common . For Tanks and Temples, following NoPe-NeRF bian2023nope and CF-3DGS fu2023colmap , we assess both novel view synthesis quality and pose estimation accuracy across 8 scenes, spanning both indoor and outdoor environments. In each case, we use 7 out of every 8 frames from the video clips as training data and evaluate the novel view synthesis on the remaining frame except Family. For CO3D-V2, containing thousands of object-centric videos where the camera orbits the object, recovering camera poses is significantly more challenging due to the complex and large camera motions. We follow the experimental settings of CF-3DGS fu2023colmap to select the same 10 scenes from different object categories and apply the same procedure to divide training and testing sets.
@@
NeRF-mm wang2021nerf - $\sim$13h33min 0.5313 20.02 0.5450 0.035
BARF lin2021barf GT K $\sim$20h 0.6075 23.42 0.5362 0.078
NoPe-NeRF bian2023nope GT K $\sim$30h 0.7125 25.49 0.4113 0.013
CF-3DGS fu2023colmap GT K $\sim$2h20min 0.9213 31.14 0.0859 0.004
@@
Metrics: We assess our approach on two key tasks: generating novel views and estimating camera poses. For the task of novel view synthesis, we evaluate performance using common metrics such as PSNR, SSIM wang2004image , and LPIPS zhang2018unreasonable . In terms of camera pose estimation, we evaluate Absolute Trajectory Error (ATE) sturm2012benchmark and utilize COLMAP-generated poses from all dense views as our ground-truth. While Relative Pose Error (RPE) sturm2012benchmark evaluates the local consistency of relative transformations between consecutive frames, it can be sensitive to discrepancies in intrinsic parameters such as focal length. ATE provides a more comprehensive measure of global trajectory accuracy and is better aligned with the goals of our method, which emphasizes globally consistent 3D reconstruction sturm2012benchmark . As such, we prioritize ATE as the primary metric for evaluating the poses of VideoLifter.
@@ 4.2 Quantitative Evaluations
To quantitatively evaluate the quality of synthesized novel views, we present the results in Tab. 3 for the Tanks and Temples dataset and Tab. 4 for the CO3D-V2 dataset. Baseline models were re-trained using their officially released code to ensure a fair comparison of training time. Compared to other self-calibrating radiance field methods, our approach achieves superior performance in terms of efficiency and rendering quality, which is largely thanks to our decoupled fragment registration and hierarchical alignment process. Compared to the most relevant baseline CF-3DGS fu2023colmap , we reduce $>$80% training time yet get $>$0.012 LPIPS improvement on Tanks and Temples, and reduce $>$85% training time yet get $>$0.12 LPIPS improvement on CO3D-V2 dataset. Note that our VideoLifter does not require any ground-truth camera parameters, making it more adaptable to scenes that do not have or fail to get precomputed intrinsics from COLMAP. Compared to NeRFmm wang2021nerf , which also does not need ground-truth camera parameters, our VideoLifter delivers much better quality and much less training time. Detailed per-scene breakdown results could be found in the Supplementary.
@@
NeRF-mm wang2021nerf - $\sim$17h22min 0.4380 13.43 0.7058 0.061
NoPe-NeRF bian2023nope GT K $\sim$35h 0.7030 25.54 0.5190 0.055
CF-3DGS fu2023colmap GT K $\sim$2h55min 0.6821 22.98 0.3515 0.014
@@ 4.3 Qualitative Evaluations
As shown in Fig. 6, for large-scale scenes from the Tanks and Temples dataset, thanks to the hierarchical design in VideoLifter, our method consistently produces sharper details among all test views, and preserves fine details that are well-optimized within each fragment. For the CO3D-V2 dataset, which includes 360-degree scenes with complex trajectories, achieving a globally consistent 3D reconstruction without any COLMAP initialization is even more challenging. Baselines that rely on monocular depth prediction to unproject images into point clouds often suffer from depth scale inconsistencies, making them fragile and prone to failure. Even CF-3DGS, which uses the more robust ZoeDepth monocular depth estimator bhat2023zoedepth , encounters severe failures on CO3D-V2. In contrast, VideoLifter leverages 3D geometry priors to achieve robust registration, making it highly adaptable and resilient in challenging settings. More results could be found in the Supplementary.
4.4 Ablation Study
Local Fragment Registration. Replacing it with direct MASt3R multi-view stereo initialization
@@ 4.4 Ablation Study
Local: Use MASt3R MVS Init. fan2024instantsplat 35min 0.8582 27.91 0.1768
@@
increases training time and lowers reconstruction quality, suggesting that outputs directly from MASt3R are less accurate and introduce errors to Gaussian optimization, especially in the challenging video setup.
Hierarchical Gaussian Alignment. Removing our hierarchical alignment and instead adding local Gaussians sequentially (as in CF-3DGS fu2023colmap ) prolongs training and hurts performance, showing the efficiency and accuracy gains from our hierarchical design.
@@
Confidence Threshold $\beta$. Setting $\beta$ too low allows fewer Gaussians to merge, leading to under-reconstructed areas, while a high $\beta$ merges too many Gaussians, slowing down training.
5 Conclusion and Limitations
We presented VideoLifter, a framework for efficient 3D scene reconstruction from monocular videos without relying on pre-computed camera poses or known intrinsics. By leveraging learning-based stereo priors and a hierarchical alignment strategy with 3D Gaussian splatting, VideoLifter produces dense, globally consistent reconstructions with significantly reduced computational overhead compared to prior methods fu2023colmap ; bian2023nope . A key limitation, shared with prior pose-free methods (e.g., CF-3DGS fu2023colmap ), is the assumption of a pinhole camera model. Extending to more general camera models remains an important direction for future work.